Apple Daney SPN Apple Bing Yahoo Geology 1100 nds 1Quiet NYC Onine P..ing Payments 5. The daia in columes 1 and 2 in the table beliow ane for a prvate cosed economy Instructions: For all pas, enter your anewers as whole numeen Eyou are nis economy to Intemational tde by Including the export and import Spures of columns 3 and4 F in the gray shaded cels in columns 5 and 6 Determine the equlibnu GOP tor the open economy What is the change in equlibrium GDP caused by the adsion of net expors? hat would be net eports and ne eoanncopnpons were S10-greaw at each ?vel of 00P?Filinthe gay shaded 8
media%2Fdd3%2Fdd3729ea-7d52-4468-9fda-c3
Answer #1

(a) Equilibrium level of GDP is when real domestic output is equal to aggregate expenditure in private closed economy . And it is equal ,when GDP level is $400 billion.Hence, equilibrium level of GDP in private closed economy = $400 billion.

(b)

(1) (2) (3) (4) (5) (6)
Real domestic output (Billions) Aggregate expenditure, private closed economy (Billions) Exports (Billions) Imports (Billions) Net exports (Billions) Aggregate expenditures, private open economy (Billions)
200 240 20 30 (20-30) =-10 (240-10)= 230
250 280 20 30 -10 (280-10)=270
300 320 20 30 -10 (320-10)=310
350 360 20 30 -10 (360-10)=350
400 400 20 30 -10 (400-10)=390
450 440 20 30 -10 (440-10)=430
500 480 20 30 -10 (480-10)=470
550 520 20 30 -10 (520-10)=510

Now, equilibrium level of GDP for the open economy is when Real domestic output is equal to aggregate expenditure in open economy . Hence , equilibrium level of GDP= $ 350 billion.

Change in equilibrium level of GDP caused by addition of net exports = $(400-350) billion = $50 billion.

(c) Now, if imports are $10 billion greater at each level of GDP, we get :

(1) (2) (3) (4) (5) (6)
Real domestic output (Billions) Aggregate expenditure, private closed economy (Billions) Exports (Billions) Imports (Billions) Net exports (Billions) Aggregate expenditures, private open economy (Billions)
200 240 20 40 (20-40) =-20 (240-20)= 220
250 280 20 30 -20 (280-20)=260
300 320 20 30 -20 (320-20)=300
350 360 20 30 -20 (360-20)=340
400 400 20 30 -20 (400-20)=380
450 440 20 30 -20 (440-20)=420
500 480 20 30 -20 (480-20)=460
550 520 20 30 -20 (520-10)=500

Now, equilibrium level of GDP for the open economy is $300 billion.